At an amusement park there is a ride in which cylindrically shaped chambers spin around a central axis. People sit in seats facing the axis, their backs against the outer wall. At one instant the outer wall moves at a speed of 2.60 m/s, and an 66.4-kg person feels a 189-N force pressing against his back. What is the radius of a chamber?
1
Expert's answer
2020-11-23T10:32:12-0500
"F=\frac{mv^2}{r}\\189=\frac{(66.4)(2.6)^2}{r}\\r=2.37\ m"